OpenStudy (reynad):

Helium is pumped into a spherical balloon at a rate of 25m^3/min. At what rate is the radius of the balloon changing when the diameter is 4m?

OpenStudy (campbell_st):

well it looks like a related rates question the volume of the sphere is \[V = \frac{4}{3} \pi r^3\] find dV/dr, the rate of change in the volume with respect to the radius you know \[\frac{dV}{dt} = 25 \] and you need to find \[\frac{dr}{dt}\] so using related rates \[\frac{dV}{dt} = \frac{dV}{dr} \times \frac{dr}{dt}\] make \[\frac{dr}{dt} ~~the ~~subject\] then substitute r = 2 hope it helps